# Neural Posterior Estimation for Cataloging Astronomical Images from the Legacy Survey of Space and Time

### v1

@@ 5.4 Galaxy shape estimation
Observed galaxy shapes provide a measurable signature of cosmic shear (D. J. Bacon et al., 2000). Measurement of many galaxies’ shapes enables reconstructing mass maps that trace the large-scale dark matter distribution (N. Jeffrey et al., 2021), constraining cosmological parameters such as $S_{8}$ (M. Asgari et al., 2021), and inferring connections between galaxies and their dark matter halos (J. Clampitt et al., 2017).
@@ 6.2 The potential of NPE to approximate the posteriors
Credible intervals based on our posterior approximations show our NPE approximations to be well calibrated, although slightly (several percent) too large. Overdispersion is not necessarily a problem, as we typically do not expect perfect calibration from variational inference, which has restrictions in the form of the variational distribution that are not shared by the exact posterior. In particular, our variational distribution is composed of simple parametric distributions. Because NPE minimizes the forward (a.k.a. exclusive) KL divergence, overdispersed approximations often result when the exact posterior cannot be recovered due to restrictions on the form of the approximation (C. Bishop, 2006, Ch. 10).
If this overdispersion were problematic, our variational distribution could be made more flexible. For example, we could use mixture models (C. M. Bishop, 1994) or normalizing flows (G. Papamakarios et al., 2021) to define our variational distribution. The ease with which we can enhance the flexibility of our variational distribution used by NPE is not shared by traditional (ELBO-based) approaches to variational inference, which typically require the reparameterization trick or, for discrete variables, continuous relaxation. In contrast, NPE imposes almost no restrictions on the form of the variational distribution.
However, while we have a clear path to decreasing miscalibration through the introduction of successively more flexible variational distributions, this path likely will not converge to perfect calibration. Sensitivity analyses will be needed to assess whether the miscalibration is substantive for any particular task. Such analyses should take into account how catalogs are used for the task in question. It often will not suffice to assess metrics that are averages across light sources, as we have done in this article, as many uses of catalogs instead involve population-level analysis (C. M. Schafer, 2015). In population-level analysis, slight miscalibration in source parameters (e.g., galaxy ellipticity) can compound, leading to large errors in the inference of population-level quantities.
@@ 6.3 Potential model misspecification
NPE can be particularly vulnerable to model misspecification due to its reliance on simulated data. If the simulator is misspecified due to uncertainty about the true data generating process, the resulting inference network might be applied to out-of-distribution settings, necessitating extrapolation by the neural network. In this setting, typical neural networks are known to perform poorly (X. Shen & N. Meinshausen, 2024).
Fortunately, several domain adaptation techniques can mitigate these issues. One method, proposed by D. Ward et al. (2022), introduces a simple parametric error model for the observed data, which is then stochastically inverted. Another method, by M. Gloeckler et al. (2023), involves regularizing the Fisher information matrix of the model parameter with respect to the data, inspired by considerations of adversarial attacks. Further, D. Huang et al. (2023) suggested adding a regularizer to the NPE objective based on the maximum mean difference (MMD). A further technique by A. Wehenkel et al. (2024) employs a small calibration set consisting of pairs of real-world observations and ground truth, which for our work could potentially come from more precise but less available instruments that produce smaller volumes of data (e.g., spectrographs).
@@ 6.4 A path forward: nonparametric modeling
Best practice for a Bayesian workflow involves iterative development guided by predictive checks of models with varying complexities (A. Gelman et al., 2020). This iterative process allows analysts to progressively refine their models, ultimately addressing various sources of model misspecification. By incrementally adding uncertainty to a simulator based on predictive checks and domain expertise, we can reduce the degree of model misspecification.
@@
There has been work developing deep generative models of galaxy images, and indeed, the models may be less misspecified (J. Regier et al., 2015; F. Lanusse et al., 2021; M. J. Smith et al., 2022). However, it is unclear how to train these models as training requires access to segmented galaxy images, which cannot be reliably found without an accurate galaxy model. In contrast, AstroPHOT (C. J. Stone et al., 2023) offers a more flexible model, specifically designed to fit images. AstroPHOT provides nonparametric component models, including Fourier- and spline-based object models, composite models, and extensible modules. This flexibility may make AstroPHOT a more appropriate simulator than GalSim for the future development of our inference method.
@@
The expectation-maximization (EM) algorithm could provide an alternative to explicitly setting a hierarchical prior. Although EM is not globally convergent, it improves the marginal likelihood of the data with each iteration. Further, for identifiable model parameters (or identifiability transformations of model parameters), with strong regularity conditions, the EM algorithm can be a consistent estimation procedure (S. Balakrishnan et al., 2017).
NPE can approximate the E-step in an EM algorithm. In this approach, NPE infers posteriors for the local random variables, while the M-step updates the model parameters based on the current posterior approximation. Whether this procedure inherits the desirable properties of EM depends on the accuracy of the approximation learned by NPE. The global convergence properties of NPE (D. McNamara et al., 2024) and the flexibility of modern variational approximations, such as normalizing flows and conditional diffusion models (L. Sharrock et al., 2024), enable NPE to solve the E-step almost exactly, which is important for retaining the theoretical properties of the EM algorithm. The proposed work can hence be viewed as a powerful approximate solution to the E-step of the EM algorithm, which, with conditions, can be a consistent estimation procedure.
@@
This material is based on work supported by the National Science Foundation under Grant No. 2209720 and the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0023714. This paper has undergone internal review in the LSST Dark Energy Science Collaboration. The internal reviewers were Chad Schafer and Grant Merz. The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy and the LSST Discovery Alliance in the United States. DESC uses resources of the IN2P3 Computing Center ($C$-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515.
@@ Appendix A Interpreting DC2 catalogs
B. Abolfathi et al. (2021b) provides a comprehensive guide to the DC2 dataset. Here, we identify and describe just the subset of DC2 fields used in our comparisons.
@@
Galaxy Shape Ixx_pixel Truth Match - second moment (xx) of source intensity
Iyy_pixel - second moment (yy) of source intensity
@@
PSF IxxPSF_pixel_<band> Truth Match pixel2 second moment (xx) of the PSF in a band
@@
IyyPSD_pixel_<band> pixel2 second moment (yy) of the PSF in a band

### v2

@@ 5.4 Galaxy shape estimation
Observed galaxy shapes provide a measurable signature of cosmic shear (Bacon et al., 2000). Measurement of many galaxies’ shapes enables reconstructing mass maps that trace the large-scale dark matter distribution (Jeffrey et al., 2021), constraining cosmological parameters such as $S_{8}$ (Asgari et al., 2021), and inferring connections between galaxies and their dark matter halos (Clampitt et al., 2017).
@@ 6.2 The potential of NPE to approximate the posteriors
Credible intervals based on our posterior approximations show our NPE approximations to be well calibrated, although slightly (several percent) too large. Overdispersion is not necessarily a problem, as we typically do not expect perfect calibration from variational inference, which has restrictions in the form of the variational distribution that are not shared by the exact posterior. In particular, our variational distribution is composed of simple parametric distributions. Because NPE minimizes the forward (a.k.a. exclusive) KL divergence, overdispersed approximations often result when the exact posterior cannot be recovered due to restrictions on the form of the approximation (Bishop, 2006, Ch. 10).
If this overdispersion were problematic, our variational distribution could be made more flexible. For example, we could use mixture models (Bishop, 1994) or normalizing flows (Papamakarios et al., 2021) to define our variational distribution. The ease with which we can enhance the flexibility of our variational distribution used by NPE is not shared by traditional (ELBO-based) approaches to variational inference, which typically require the reparameterization trick or, for discrete variables, continuous relaxation. In contrast, NPE imposes almost no restrictions on the form of the variational distribution.
However, while we have a clear path to decreasing miscalibration through the introduction of successively more flexible variational distributions, this path likely will not converge to perfect calibration. Sensitivity analyses will be needed to assess whether the miscalibration is substantive for any particular task. Such analyses should take into account how catalogs are used for the task in question. It often will not suffice to assess metrics that are averages across light sources, as we have done in this article, as many uses of catalogs instead involve population-level analysis (Schafer, 2015). In population-level analysis, slight miscalibration in source parameters (e.g., galaxy ellipticity) can compound, leading to large errors in the inference of population-level quantities.
@@ 6.3 Potential model misspecification
NPE can be particularly vulnerable to model misspecification due to its reliance on simulated data. If the simulator is misspecified due to uncertainty about the true data generating process, the resulting inference network might be applied to out-of-distribution settings, necessitating extrapolation by the neural network. In this setting, typical neural networks are known to perform poorly (Shen & Meinshausen, 2024).
Fortunately, several domain adaptation techniques can mitigate these issues. One method, proposed by Ward et al. (2022), introduces a simple parametric error model for the observed data, which is then stochastically inverted. Another method, by Gloeckler et al. (2023), involves regularizing the Fisher information matrix of the model parameter with respect to the data, inspired by considerations of adversarial attacks. Further, Huang et al. (2023) suggested adding a regularizer to the NPE objective based on the maximum mean difference (MMD). A further technique by Wehenkel et al. (2024) employs a small calibration set consisting of pairs of real-world observations and ground truth, which for our work could potentially come from more precise but less available instruments that produce smaller volumes of data (e.g., spectrographs).
@@ 6.4 A path forward: nonparametric modeling
Best practice for a Bayesian workflow involves iterative development guided by predictive checks of models with varying complexities (Gelman et al., 2020). This iterative process allows analysts to progressively refine their models, ultimately addressing various sources of model misspecification. By incrementally adding uncertainty to a simulator based on predictive checks and domain expertise, we can reduce the degree of model misspecification.
@@
There has been work developing deep generative models of galaxy images, and indeed, the models may be less misspecified (Regier et al., 2015; Lanusse et al., 2021; Smith et al., 2022). However, it is unclear how to train these models as training requires access to segmented galaxy images, which cannot be reliably found without an accurate galaxy model. In contrast, AstroPHOT (Stone et al., 2023) offers a more flexible model, specifically designed to fit images. AstroPHOT provides nonparametric component models, including Fourier- and spline-based object models, composite models, and extensible modules. This flexibility may make AstroPHOT a more appropriate simulator than GalSim for the future development of our inference method.
@@
The expectation-maximization (EM) algorithm could provide an alternative to explicitly setting a hierarchical prior. Although EM is not globally convergent, it improves the marginal likelihood of the data with each iteration. Further, for identifiable model parameters (or identifiability transformations of model parameters), with strong regularity conditions, the EM algorithm can be a consistent estimation procedure (Balakrishnan et al., 2017).
NPE can approximate the E-step in an EM algorithm. In this approach, NPE infers posteriors for the local random variables, while the M-step updates the model parameters based on the current posterior approximation. Whether this procedure inherits the desirable properties of EM depends on the accuracy of the approximation learned by NPE. The global convergence properties of NPE (McNamara et al., 2024) and the flexibility of modern variational approximations, such as normalizing flows and conditional diffusion models (Sharrock et al., 2024), enable NPE to solve the E-step almost exactly, which is important for retaining the theoretical properties of the EM algorithm. The proposed work can hence be viewed as a powerful approximate solution to the E-step of the EM algorithm, which, with conditions, can be a consistent estimation procedure.
@@
Acknowledgements
This material is based on work supported by the National Science Foundation under Grant No. 2209720 and the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0023714.
This paper has undergone internal review in the LSST Dark Energy Science Collaboration. The internal reviewers were Chad Schafer and Grant Merz. The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy and the LSST Discovery Alliance in the United States. DESC uses resources of the IN2P3 Computing Center ($C$-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515.
@@ Appendix A Interpreting DC2 catalogs
Abolfathi et al. (2021b) provides a comprehensive guide to the DC2 dataset. Here, we identify and describe just the subset of DC2 fields used in our comparisons.
@@
Galaxy Shape Ixx_pixel Truth Match - second moment ($x$) of source intensity
Iyy_pixel - second moment ($y$) of source intensity
@@
PSF IxxPSF_pixel_<band> Truth Match pixel2 second moment ($x$) of the PSF in a band
@@
IyyPSD_pixel_<band> pixel2 second moment ($y$) of the PSF in a band